# Exploration and Incentivized Participation in Randomized Trials

### v11

@@ 4.2 Our mechanism: Main stage
$\mathtt{gap}(\mathcal{P}):=\textstyle\operatornamewithlimits{\mathbb{E}}_{\psi\sim\mathcal{P}}\left[\,U_{\psi}(a^{*}(\psi))-\tfrac{1}{n}\cdot\sum_{a\in[n]}U_{\psi}(a)\,\right]$
@@
$\mathbb{P}_{\mathtt{Bayes}}\left[\,U_{\psi}(\,a^{*}(\overline{\psi})\,)\geq U_{\psi}(\,a^{*}(\psi)\,)-2\delta\mid\psi^{*}=\psi\,\right]\geq 1-p_{\mathtt{err}}$
@@ 4.4 Lower bound
Thus, Theorem 4.8 allows per-round policies that change over time and can themselves be jointly chosen at random. In rounds $t\not\in S$, the policies can depend on all past observations. Importantly, this includes “multi-stage” mechanisms that partition the time into a constant number of contiguous “stages” and are non-data-adaptive within each (possibly with a data-adaptive warm-up stage of duration, say, $<T/2$). Most medical trials in practice fit this model, see Footnote 9.
@@
[…]
@@ 5.2 Model: heterogeneous agents
To extend the BIR property, we focus on Bayesian-expected utility from truthful reporting. Let $\mathcal{E}^{\mathtt{tru}}_{t}$ be the event that all agents $s\leq t-1$ have reported their private type truthfully: $\theta_{\mathtt{rep}}^{(s)}=\theta_{\mathtt{pri}}^{(s)}$. Let $\mathbb{E}_{\mathtt{Bayes}}\left[\,\cdot\mid\mathcal{E}\,\right]$ denote expectation with respect to agents’ Bayesian beliefs given event $\mathcal{E}$ [^24] We use non-standard notation for the expectation to emphasize that it only needs to be well-defined under event $\mathcal{E}$, even though it may be ill-defined unconditionally.
@@ 5.3 Our results
$\hat{\sigma}(\psi,\theta):=\sigma(\psi,\theta^{*}(\theta)),\text{where }\theta^{*}(\theta)\in\operatornamewithlimits{argmax}_{\theta^{\prime}\in\Theta}U_{\mathcal{P}}\left(\,\sigma(\,\cdot\,,\theta^{\prime}),\theta\,\right),$
@@
$U^{*}_{\mathcal{P}}(\theta)$ $:=\operatornamewithlimits{\mathbb{E}}_{\psi\sim\mathcal{P}}\left[\,U_{\psi}\left(\,a^{*}(\psi,\theta),\,\theta\,\right)\,$
@@
$U^{*}_{\mathcal{P}}(\theta,\theta_{\mathtt{rep}})$ $:=\operatornamewithlimits{\mathbb{E}}_{\psi\sim\mathcal{P}}\left[\,U_{\psi}\left(\,a^{*}(\psi,\theta^{\prime}),\,\theta\,\right)\,\right],\quad\text{where }\theta^{\prime}=(\theta_{\mathtt{pub}},\theta_{\mathtt{rep}})$
@@
We can interpret the best arm $a^{*}()=a^{*}(\psi^{*},\theta)$ as a state-aware recommendation policy, called the best-arm policy. Then condition (5.9) states that $a^{*}()$ is strictly BIR and strictly BIC.
@@
[…]
@@
Our mechanism is defined as follows. We choose uniformly between $\sigma^{\mathtt{opt}}$ and the best-arm policy $a^{*}()$. Formally: in each round $t$ of the main stage, the sampling distribution $p_{t}$ is the average between two distributions: $\sigma^{\mathtt{opt}}(\psi_{\mathtt{MLE}},\theta)$ and $a^{*}(\psi_{\mathtt{MLE}},\theta)$, where $\theta=\left(\,\theta_{\mathtt{pub}}^{(t)},\,\theta_{\mathtt{rep}}^{(t)}\,\right)$ is the observed type. For estimator $\widehat{f}_{\mathtt{mech}}$, we use the IPS estimator (3.5). This completes the description of the main stage. We denote our mechanism as $\mathtt{TwoStageMech}(T_{0},\sigma^{\mathtt{opt}},a^{*}())$.
@@
Assume (5.9). Let $\mathtt{mech}$ be the mechanism $\mathtt{TwoStageMech}(T_{0},\sigma^{\mathtt{opt}},a^{*}())$. Then:
@@
    $\mathbb{P}_{\mathtt{Bayes}}\left[\,\text{each (arm, public type) pair appears in $\geq N_{\mathcal{P}}$ rounds}\,\right]\geq 1-\delta_{\mathcal{P}},$ (5.10)
@@
Main stage: extensions. It may be desirable to use relaxed versions of the benchmark-optimizing policy $\sigma^{\mathtt{opt}}$ and the best-arm policy $a^{*}()$. Indeed, the exact versions may be too difficult to compute or too complex to implement in practice; also, the prior $\mathcal{P}$ or the utility structure may be not fully known to the mechanism. We relax the benchmark-optimizing policy by allowing additive slack $\delta>0$, which shows up in the final regret bound. We relax the best-arm policy as an arbitrary state-aware policy $\widehat{\sigma}$ which is BIR and BIC by some additive margin $\eta\leq\eta^{*}$; this margin only affects the number of warm-up samples.
@@
A recommendation policy is called $\eta$-BIR (resp., $\eta$-BIC) for some $\eta\geq 0$ if satisfies Eq. (5.4) (resp., Eq. (5.5)) with $+\eta$ added to the right-hand side.
@@
Assume (5.9). Let $\widehat{\sigma}^{\mathtt{opt}}$ be a state-aware policy which optimizes the $\inf$ in the benchmark (5.7) up to an additive factor $\delta\geq 0$. Let $\widehat{\sigma}$ be a state-aware policy which is $\eta$-BIR and $\eta$-BIC, for some $\eta\in(0,\eta^{*}]$. Then mechanism $\mathtt{TwoStageMech}(T_{0},\widehat{\sigma}^{\mathtt{opt}},\widehat{\sigma})$ satisfies part (b) in Theorem 5.10 with parameters $\delta_{\mathcal{P}}=\eta/8$ and $N_{\mathcal{P}}=N_{\mathcal{P}}(\eta)$, and satisfies
@@ 6 Heterogeneous agents with estimated type frequencies
[…]
@@ 6.1 Results
We reuse mechanism $\mathtt{TwoStageMech}(T_{0},\sigma^{\mathtt{opt}},a^{*}())$ from Section 5 where $a^{*}()$ is the best-arm policy and $\sigma^{\mathtt{opt}}=\sigma^{\mathtt{opt}}(\mathcal{P},\widetilde{F}_{\mathtt{type}})$ is the benchmark-optimizing policy relative to the estimated type frequencies. The number of warm-up samples is given by the same Eq. (5.12).
@@
Assume (5.9). Let $\widetilde{F}_{\mathtt{type}}$ be some distribution over types. Consider mechanism $\mathtt{TwoStageMech}(T_{0},\sigma^{\mathtt{opt}},a^{*}())$, where policy $\sigma^{\mathtt{opt}}=\sigma^{\mathtt{opt}}(\widetilde{F}_{\mathtt{type}})$ optimizes the benchmark for distribution $\widetilde{F}_{\mathtt{type}}$. This mechanism satisfies part (b) in Theorem 5.10 with parameters $\delta_{\mathcal{P}}=\eta^{*}/8$ and $N_{\mathcal{P}}=N_{\mathcal{P}}(\eta^{*})$, as per Eqns. (5.11) and (5.12). The performance guarantee is as follows:
@@
Theorem 6.2 can be extended to “relaxed” versions of $a^{*}()$ and $\sigma^{\mathtt{opt}}$, just like in Theorem 5.12. The exact formulation is omitted, as we believe it does not yield additional insights.
@@ 7 Conclusions and open questions
We introduce a model for incentivized participation in clinical trials. The model combines the statistical objective (which is standard for clinical trials and adversarial in nature) from the economic constraints (which are standard in economic theory and based on Bayesian-stochastic beliefs). The model emphasizes homogeneous agents as the paradigmatic case, and extends to heterogeneous agents. The latter are endowed with public types that affect objective outcomes, and private types that affect subjective preferences over these outcomes.
@@
Our work sets the stage for further investigation. First, while we mitigate the influence of rare-but-difficult agent types on the main stage (via estimated type frequencies), can one do it for the warm-up stage, too? That is, reduce its duration and/or reuse its data for the counterfactual estimates. Second, can one match the result for known type frequencies if the type frequencies are instead learned on the fly? Third, how to think about incentivized RCTs if the agents cannot fully quantify their own risk preferences, and therefore cannot fully report them?
@@ Appendix A Simple numerical examples
Homogeneous agents. There are $n=2$ treatments (henceforth, “arms”), the “old” and the “new”. As per our model, agents believe each arm gives the same utility distribution for all agents, and the state (which we identify with a mapping from arms to utility distributions) is initially drawn from some prior $\mathcal{P}$ (shared by all agents). Specifically, the old arm gives mean utility 1 with probability $\nicefrac{{2}}{{3}}$ and utility 0 with probability $\nicefrac{{1}}{{3}}$. The new arm gives utility 2 and -1 with probability $\nicefrac{{1}}{{2}}$ each. The prior is independent across arms. Given the mean utilities (and the fact that the realized utilities lie in the $[0,1]$ interval), the noise shape can be arbitrary for this example. [^27] Thus, in the formalism of our model, we have four possible states, with mean utility pairs $(1,2)$, $(1,-1)$, $(0,2)$ and $(0,-1)$, and respective prior probabilities $\nicefrac{{1}}{{3}}$, $\nicefrac{{1}}{{3}}$, $\nicefrac{{1}}{{6}}$, $\nicefrac{{1}}{{6}}$. Agents’ alternative to participation is to use the old arm, whose prior-mean utility is $\nicefrac{{2}}{{3}}$; so, the “outside option” is $U_{\mathtt{out}}=\nicefrac{{2}}{{3}}$. Note that it is ex-ante preferred to the new arm (whose prior-mean utility is only $\nicefrac{{1}}{{2}}$), so the agents are incentivized against participation if the arm probabilities are data-independent.
@@
Our participation-incentivizing RCT design from Section 4 works out as follows. Leveraging Lemma 4.2, consider $\epsilon^{*}$, the largest $\epsilon$ such that $(\epsilon,\psi^{*})\text{-$\mathtt{greedy}$}$ policy is BIR. We claim $\epsilon^{*}=\nicefrac{{8}}{{9}}$. Indeed, note that $\epsilon^{*}$ is is the unique $\epsilon$ such that the Bayesian-expected utility of $(\epsilon,\psi^{*})\text{-$\mathtt{greedy}$}$, call it $U(\epsilon)$, equals the outside option $U_{\mathtt{out}}$. And one can check that $U(\nicefrac{{8}}{{9}})=\nicefrac{{2}}{{3}}=U_{\mathtt{out}}$. (One way to do that is to sum over the four possible states, as per Footnote 27, and for each state calculate the expected utility of the policy; we omit the easy details.)
Thus, we obtain the benchmark value $\mathtt{bench}(\mathcal{P})=n/\epsilon^{*}=\nicefrac{{9}}{{4}}$.
@@
Now let us calculate $N_{\mathcal{P}}$, the sufficient number of samples for the warm-up stage, as per Eq. (4.2) in our analysis. [^28] While the expression in Eq. (4.2) has an appealing “interpretability”, characterizing $N_{\mathcal{P}}$ in terms of the prior gap $\mathtt{gap}(\mathcal{P})$ and the benchmark $\mathtt{bench}(\mathcal{P})$, we suspect one can improve over Eq. (4.2) with a more elaborate analysis. The prior-expected utility of the best arm is $U_{\mathcal{P}}(a^{*}())=\nicefrac{{4}}{{3}}$ (the calculation is similar to that for $U(\epsilon^{*})$.) So, the “prior gap” (see Eq. (4.2)) is $\mathtt{gap}(\mathcal{P})=\nicefrac{{4}}{{3}}-\nicefrac{{5}}{{12}}=\nicefrac{{11}}{{12}}$. Therefore, according to Eq. 4.3, we have $\alpha=\nicefrac{{22}}{{27}}$ and $N_{P}<144$. In other words, a warm-up stage with $144$ independent samples of each arm suffices to guarantee incentive-compatibility for the main stage. Recall that these samples may also be obtained exogenously, e.g., via paid volunteers.
@@
By assumption, with probability at least $81-\eta/$ 8, the mechanism collects at least $N_{\mathcal{P}}(\eta)$ samples for each (action, public type) pair. Denote this event by $\mathcal{E}_{0}$. For any $\psi^{\prime}\in\mathtt{support}(\mathcal{P})\setminus\left\{\,\psi^{*}\,\right\}$, there exists at least one (action, public type) pair $(a,\theta_{\mathtt{pub}})$ such that $\psi^{*}(a,\theta_{\mathtt{pub}})\neq\psi^{\prime}(a,\theta_{\mathtt{pub}})$. Therefore, under event $\mathcal{E}_{0}$ we have $|S\left(\,\psi^{*},\psi^{\prime}\,\right)|\geq N_{\mathcal{P}}(\eta)$. It follows that with probability at least $41-\eta/$ 4, event $\mathcal{E}\left(\,\psi^{*},\psi^{\prime}\,\right)$ holds for all states $\psi^{\prime}\in\mathtt{support}(\mathcal{P})\setminus\left\{\,\psi^{*}\,\right\}$. This implies Eq. (B.2), and therefore completes the proof of Claim 5.16.
@@
Fix $\tau\in[T_{0}]$. Since $\mathbb{E}_{\mathtt{Bayes}}\left[\,\hat{X}_{j}\mid\hat{X}_{j-1}\,\right]=0$ for all $j\in[T_{0}]$, the sum $\hat{X}_{[\tau]}:=\sum_{j\leq\tau}\hat{X}_{j}$ is a martingale (in the probability space induced by the agents’ beliefs). This martingale’s increments are bounded as $\hat{X}_{j}\leq\ell_{\max}(\psi,\psi^{\prime})$. Applying the standard Azuma-Hoeffding inequality with deviation term $-\tau\cdot{\rm KL}_{\min}(\psi,\psi^{\prime})$, we obtain

### v12

@@ 4.2 Our mechanism: Main stage
$\mathtt{gap}(\mathcal{P}):=\textstyle\operatornamewithlimits{\mathbb{E}}_{\psi\sim\mathcal{P}}\left[\,U_{\psi}(a^{*}(\psi))-\tfrac{1}{n}\cdot\sum_{a\in[n]}U_{\psi}(a)\,\right],$ (4.4)
@@
$\mathbb{P}_{\mathtt{Bayes}}\left[\,U_{\psi}(\,a^{*}(\overline{\psi})\,)\geq U_{\psi}(\,a^{*}(\psi)\,)-2\delta\mid\psi^{*}=\psi\,\right]\geq 1-p_{\mathtt{err}}.$ (4.6)
@@
$U_{\mathcal{P}}\left(\,(\epsilon,\overline{\psi})\text{-$\mathtt{greedy}$}\,\right)$ $=\operatornamewithlimits{\mathbb{E}}_{\psi\sim\mathcal{P}}\left[\,(1-\epsilon)\cdot U_{\psi}(a_{\overline{\psi}})+\frac{\epsilon}{n}\cdot{\sum_{a\in[n]}}U_{\psi}(a)\,\right]$
$\geq\operatornamewithlimits{\mathbb{E}}_{\psi\sim\mathcal{P}}\left[\,(1-\epsilon)\cdot U_{\psi}(a^{*}(\psi))+\frac{\epsilon}{n}\cdot{\sum_{a\in[n]}}U_{\psi}(a)\,\right]-p_{\mathtt{err}}-2\delta$
$=\operatornamewithlimits{\mathbb{E}}_{\psi\sim\mathcal{P}}\left[\,(1-2\epsilon)\cdot U_{\psi}(a^{*}(\psi))+\frac{2\epsilon}{n}\cdot{\sum_{a\in[n]}}U_{\psi}(a)\,\right]+\epsilon\cdot\mathtt{gap}(\mathcal{P})-p_{\mathtt{err}}-2\delta$
$\geq U_{\mathcal{P}}(\sigma^{*}),$
@@ 4.4 Lower bound
Thus, Theorem 4.8 allows per-round policies that change over time and can themselves be jointly chosen at random. In rounds $t\not\in S$, the policies can depend on all past observations. Importantly, this includes “multi-stage” mechanisms that partition the time into a constant number of contiguous “stages” and are non-data-adaptive within each (possibly with a data-adaptive warm-up stage of duration, say, <T/2<T/2). Most medical trials in practice fit this model, see Footnote 11.
@@
$\mathtt{ERR}\left(\,\mathtt{mech}\mid\mathtt{adv}^{\dagger}\,\right)\geq\Omega\left(\,\frac{1}{|S|^{2}}\;\max_{a\in[n]}\;\sum_{t\in S}\min\left\{\,1/\operatornamewithlimits{\mathbb{E}}\left[\,p_{t}(a)\,\right],\sqrt{|S|}\,\right\}\,\right),$ (4.8)
@@ 5.2 Model: heterogeneous agents
To extend the BIR property, we focus on Bayesian-expected utility from truthful reporting. Let $\mathcal{E}^{\mathtt{tru}}_{t}$ be the event that all agents $s\leq t-1$ have reported their private type truthfully: $\theta_{\mathtt{rep}}^{(s)}=\theta_{\mathtt{pri}}^{(s)}$. Let $\mathbb{E}_{\mathtt{Bayes}}\left[\,\cdot\mid\mathcal{E}\,\right]$ denote expectation with respect to agents’ Bayesian beliefs given event $\mathcal{E}$. [^25] We use non-standard notation for the expectation to emphasize that it only needs to be well-defined under event $\mathcal{E}$, even though it may be ill-defined unconditionally.
@@ 5.3 Our results
[…]
@@
$U^{*}_{\mathcal{P}}(\theta)$ $:=\operatornamewithlimits{\mathbb{E}}_{\psi\sim\mathcal{P}}\left[\,U_{\psi}\left(\,a^{*}(\psi,\theta),\,\theta\,\right)\,\right]$ (5.9)
@@
$U^{*}_{\mathcal{P}}(\theta,\theta_{\mathtt{rep}})$ $:=\operatornamewithlimits{\mathbb{E}}_{\psi\sim\mathcal{P}}\left[\,U_{\psi}\left(\,a^{*}(\psi,\theta^{\prime}),\,\theta\,\right)\,\right],\quad\text{where }\theta^{\prime}=(\theta_{\mathtt{pub}},\theta_{\mathtt{rep}}).$ (5.10)
@@
We can interpret the best arm $a^{*}=a^{*}(\psi^{*},\theta)$ as a state-aware recommendation policy, called the best-arm policy. Then condition (5.11) states that $a^{*}$ is strictly BIR and strictly BIC.
@@
By (5.11), the $(\epsilon,\psi^{*})\text{-$\mathtt{greedy}$}$ policy is BIC and BIR for some $\epsilon>0$. [^26] As in Section 4, $(\epsilon,\psi^{*})\text{-$\mathtt{greedy}$}$ chooses the best-arm policy in the exploitation branch. This policy satisfies the conditions under the $\inf$ in (5.7), and has a finite “benchmark value”, namely $\tfrac{1}{\epsilon}$. ∎
@@
Our mechanism is defined as follows. We choose uniformly between $\sigma^{\mathtt{opt}}$ and the best-arm policy $a^{*}$. Formally: in each round $t$ of the main stage, the sampling distribution $p_{t}$ is the average between two distributions: $\sigma^{\mathtt{opt}}(\psi_{\mathtt{MLE}},\theta)$ and $a^{*}(\psi_{\mathtt{MLE}},\theta)$, where $\theta=\left(\,\theta_{\mathtt{pub}}^{(t)},\,\theta_{\mathtt{rep}}^{(t)}\,\right)$ is the observed type. For estimator $\widehat{f}_{\mathtt{mech}}$, we use the IPS estimator (3.5). This completes the description of the main stage. We denote our mechanism as $\mathtt{TwoStageMech}(T_{0},\sigma^{\mathtt{opt}},a^{*})$.
@@
Assume (5.11). Let $\mathtt{mech}$ be the mechanism $\mathtt{TwoStageMech}(T_{0},\sigma^{\mathtt{opt}},a^{*})$. Then:
@@
[…]
@@
Main stage: extensions. It may be desirable to use relaxed versions of the benchmark-optimizing policy $\sigma^{\mathtt{opt}}$ and the best-arm policy $a^{*}$. Indeed, the exact versions may be too difficult to compute or too complex to implement in practice; also, the prior $\mathcal{P}$ or the utility structure may be not fully known to the mechanism. We relax the benchmark-optimizing policy by allowing additive slack $\delta>0$, which shows up in the final regret bound. We relax the best-arm policy as an arbitrary state-aware policy $\widehat{\sigma}$ which is BIR and BIC by some additive margin $\eta\leq\eta^{*}$; this margin only affects the number of warm-up samples.
@@
A recommendation policy is called $\eta$-BIR (resp., $\eta$-BIC) for some $\eta\geq 0$ if satisfies Eq. (5.4) (resp., Eq. (5.5)) with $\eta$ added to the right-hand side.
@@
[…]
@@ 6 Heterogeneous agents with estimated type frequencies
$F_{\mathtt{type}}(\theta\mid\mathtt{adv})=\frac{1}{T-T_{0}}\,\sum_{t>T_{0}}\;{\bf 1}_{\left\{\,\theta_{t}=\theta\,\right\}}\qquad\forall\theta\in\Theta.$ (6.1)
@@ 6.1 Results
We reuse mechanism $\mathtt{TwoStageMech}(T_{0},\sigma^{\mathtt{opt}},a^{*})$ from Section 5 where $a^{*}$ is the best-arm policy and $\sigma^{\mathtt{opt}}=\sigma^{\mathtt{opt}}(\mathcal{P},\widetilde{F}_{\mathtt{type}})$ is the benchmark-optimizing policy relative to the estimated type frequencies. The number of warm-up samples is given by the same Eq. (5.14).
@@
Assume (5.11). Let $\widetilde{F}_{\mathtt{type}}$ be some distribution over types. Consider mechanism $\mathtt{TwoStageMech}(T_{0},\sigma^{\mathtt{opt}},a^{*})$, where policy $\sigma^{\mathtt{opt}}=\sigma^{\mathtt{opt}}(\widetilde{F}_{\mathtt{type}})$ optimizes the benchmark for distribution $\widetilde{F}_{\mathtt{type}}$. This mechanism satisfies part (b) in Theorem 5.10 with parameters $\delta_{\mathcal{P}}=\eta^{*}/8$ and $N_{\mathcal{P}}=N_{\mathcal{P}}(\eta^{*})$, as per Eqns. (5.13) and (5.14). The performance guarantee is as follows:
@@
Theorem 6.2 can be extended to “relaxed” versions of $a^{*}$ and $\sigma^{\mathtt{opt}}$, just like in Theorem 5.12. The exact formulation is omitted, as we believe it does not yield additional insights.
@@ 7 Conclusions and open questions
We introduce a model for incentivized participation in RCTs. The model combines the statistical objective (which is standard for clinical trials and adversarial in nature) from the economic constraints (which are standard in economic theory and based on Bayesian-stochastic beliefs). The model emphasizes homogeneous agents as the paradigmatic case, and extends to heterogeneous agents. The latter are endowed with public types that affect objective outcomes, and private types that affect subjective preferences over these outcomes.
@@
Our work sets the stage for further investigation. First, can we reduce the sufficient duration of the warm-up stage by integrating some of the data-collection therein into the main stage? Second, while we mitigate the influence of rare-but-difficult agent types for the main stage (via estimated type frequencies), can one do it for (the sufficient duration of) the warm-up stage, too? Third, can one match our result for estimated type frequencies if they not given in advance but are instead learned on the fly? Fourth, can one meaningfully relax our “statistical model”? That is, constrain the adversarial outcomes while still retaining the need for independent per-round randomization. The goal would be to derive stronger guarantees, positive and/or negative. Fifth, and more speculatively: how to think about incentivized RCTs if the agents cannot fully quantify their own risk preferences, and therefore cannot fully report them?
@@ Appendix A Simple numerical examples
Homogeneous agents. There are $n=2$ treatments (henceforth, “arms”), the “old” and the “new”. As per our model, agents believe each arm gives the same utility distribution for all agents, and the state (which we identify with a mapping from arms to utility distributions) is initially drawn from some prior $\mathcal{P}$ (shared by all agents). Specifically, the old arm gives mean utility 1 with probability $\nicefrac{{2}}{{3}}$ and utility 0 with probability $\nicefrac{{1}}{{3}}$. The new arm gives utility 2 and -1 with probability $\nicefrac{{1}}{{2}}$ each. The prior is independent across arms. Given the mean utilities (and the fact that the realized utilities lie in the $[0,1]$ interval), the noise shape can be arbitrary for this example. [^28] Thus, in the formalism of our model, we have four possible states, with mean utility pairs $(1,2)$, (1,1)(1,-1), $(0,2)$ and (0,1)(0,-1), and respective prior probabilities $\nicefrac{{1}}{{3}}$, $\nicefrac{{1}}{{3}}$, $\nicefrac{{1}}{{6}}$, $\nicefrac{{1}}{{6}}$. Agents’ alternative to participation is to use the old arm, whose prior-mean utility is $\nicefrac{{2}}{{3}}$; so, the “outside option” is $U_{\mathtt{out}}=\nicefrac{{2}}{{3}}$. Note that it is ex-ante preferred to the new arm (whose prior-mean utility is only $\nicefrac{{1}}{{2}}$), so the agents are incentivized against participation if the arm probabilities are data-independent.
@@
Our participation-incentivizing RCT design from Section 4 works out as follows. Leveraging Lemma 4.2, consider $\epsilon^{*}$, the largest $\epsilon$ such that $(\epsilon,\psi^{*})\text{-$\mathtt{greedy}$}$ policy is BIR. We claim $\epsilon^{*}=\nicefrac{{8}}{{9}}$. Indeed, note that $\epsilon^{*}$ is is the unique $\epsilon$ such that the Bayesian-expected utility of $(\epsilon,\psi^{*})\text{-$\mathtt{greedy}$}$, call it $U(\epsilon)$, equals the outside option $U_{\mathtt{out}}$. And one can check that $U(\nicefrac{{8}}{{9}})=\nicefrac{{2}}{{3}}=U_{\mathtt{out}}$. (One way to do that is to sum over the four possible states, as per Footnote 28, and for each state calculate the expected utility of the policy; we omit the easy details.)
Thus, we obtain the benchmark value:
$\mathtt{bench}(\mathcal{P})=n/\epsilon^{*}=\nicefrac{{9}}{{4}}.$
@@
Now let us calculate $N_{\mathcal{P}}$, the sufficient number of samples for the warm-up stage, as per Eq. (4.4). The expression in Eq. (4.4) has an appealing “interpretability”, characterizing $N_{\mathcal{P}}$ in terms of the prior gap $\mathtt{gap}(\mathcal{P})$ and the benchmark $\mathtt{bench}(\mathcal{P})$. However, as stated in the Introduction, we suspect one can improve over Eq. (4.4) with a more elaborate analysis.
The calculation is as follows. The prior-expected utility of the best arm is $U_{\mathcal{P}}(a^{*})=\nicefrac{{4}}{{3}}$ (the calculation is similar to that for $U(\epsilon^{*})$.) So, the “prior gap” (see Eq. (4.4)) is $\mathtt{gap}(\mathcal{P})=\nicefrac{{4}}{{3}}-\nicefrac{{5}}{{12}}=\nicefrac{{11}}{{12}}$. Therefore, according to Eq. 4.5, we have $\alpha=\nicefrac{{22}}{{27}}$ and
$N_{P}<144.$
In other words, a warm-up stage with $144$ independent samples of each arm suffices to guarantee incentive-compatibility for the main stage. Recall that these samples may also be obtained exogenously, e.g., via paid volunteers.
@@
By assumption, with probability at least $1-\eta/8$, the mechanism collects at least $N_{\mathcal{P}}(\eta)$ samples for each (action, public type) pair. Denote this event by $\mathcal{E}_{0}$. For any $\psi^{\prime}\in\mathtt{support}(\mathcal{P})\setminus\left\{\,\psi^{*}\,\right\}$, there exists at least one (action, public type) pair $(a,\theta_{\mathtt{pub}})$ such that $\psi^{*}(a,\theta_{\mathtt{pub}})\neq\psi^{\prime}(a,\theta_{\mathtt{pub}})$. Therefore, under event $\mathcal{E}_{0}$ we have $|S\left(\,\psi^{*},\psi^{\prime}\,\right)|\geq N_{\mathcal{P}}(\eta)$. It follows that with probability at least $1-\eta/4$, event $\mathcal{E}\left(\,\psi^{*},\psi^{\prime}\,\right)$ holds for all states $\psi^{\prime}\in\mathtt{support}(\mathcal{P})\setminus\left\{\,\psi^{*}\,\right\}$. This implies Eq. (B.2), and therefore completes the proof of Claim 5.16.
@@
Fix $\tau\in[T_{0}]$. Since $\mathbb{E}_{\mathtt{Bayes}}\left[\,\hat{X}_{j}\mid\hat{X}_{j-1}\,\right]=0$ for all $j\in[T_{0}]$, the sum $\hat{X}_{[\tau]}:=\sum_{j\leq\tau}\hat{X}_{j}$ is a martingale (in the probability space induced by the agents’ beliefs). This martingale’s increments are bounded as $\hat{X}_{j}\leq\ell_{\max}(\psi,\psi^{\prime})$. Applying the standard Azuma-Hoeffding inequality with deviation term $\tau\cdot{\rm KL}_{\min}(\psi,\psi^{\prime})$, we obtain